OpenStudy (anonymous):

1. Determine whether the equation has one solution, no solution, or an infinite number of solutions. 3(x + 5) + x = 4x

OpenStudy (anonymous):

@iGreen.

Directrix (directrix):

Solve this for x: 3(x + 5) + x = 4x

OpenStudy (anonymous):

it would be 3x+15+x=4x and that would be 4x+15=4x @Directrix

Directrix (directrix):

Just a sec

Directrix (directrix):

>> 4x+15=4x That is correct. What do you get when you subtract 4x from both sides?

Directrix (directrix):